# CHaRNet: Conditioned Heatmap Regression for Robust Dental Landmark Localization

### v4

@@ 5 Experiments
In this section, we validate the effectiveness of our CHaR module for the challenging task of 3D landmark detection in dentition models, particularly in cases with missing teeth. We demonstrate that integrating the CHaR module into various point cloud learning models significantly improves localization performance. To ensure robustness, we evaluated across five point cloud processing architectures spanning different families, comparing baseline models with their CHaR-augmented versions.
A description of the used evaluation metrics is introduced in Section 5.1, followed by a detailed presentation of the results in Section 5.2.
5.1 Evaluation metrics
In order to evaluate the performance of the tooth landmark detection and tooth classification in 3D dental models, we have used the mean Euclidean distance error (MEDE) and the mean success ratio (MSR) for the landmark detection task, and F1 score to also evaluate the classification performance.
Let $N$ be the number of landmarks in a dental model. Then, for each landmark $i\in[1,N]$, we can define the Euclidean distance $D_{i}$ as:
$D_{i}=\sqrt{\sum_{j=1}^{3}(y_{ij}-\hat{y}_{ij})^{2}}$ (7)
where $y_{i}$ and $\hat{y}_{i}$ are the actual and predicted landmarks, respectively. Subsequently, we define the MEDE as follows.
$MEDE=\frac{\sum_{i=1}^{N}{D_{i}}}{N}$ (8)
Finally, the MSR is expressed as
$MSR=\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\times 100\%,\quad\delta_{i}=\begin{cases}1&\text{if }D_{i}\leq r,\\ 0&\text{if }D_{i}>r\end{cases}$ (9)
where $r$ is the radius in $mm$ of the sphere that delimits the space for a predicted landmark to be considered a success. In our case, we define $r=1mm$.
Finally, since our CHaR module makes use of the teeth classification to condition the final landmark location, we also considered the F1 score to evaluate the performance in this regard:
$F1_{t}=\frac{2*TP_{t}}{2*TP_{t}+FP_{t}+FN_{t}}$ (10)
where $TP_{t}$, $FP_{t}$, $FN_{t}$ are true positives (TP), false positives (FP), and false negatives (FN) for each tooth $t$. To compute the overall or macro-averaged, we take the mean across all teeth:
$F1_{macro}=\frac{1}{T}\sum_{t=1}^{T}F1_{t}$ (11)
where $T=16$ is the total number of teeth.
5.2 Results
The quantitative 3D landmark detection results for the five point cloud base models, namely PointNet++, DGCNN, PointTransformer, PointNeXt, and PointMLP with and without CHaR, are compared in Table 5 and Table 6, using MEDE and MSR as evaluation metrics, respectively. Three key observations can be made. First, the CHaR-augmented versions consistently outperform their base counterparts, demonstrating the effectiveness of the proposed method. Second, the models perform well on typical dental models, whereas cases involving missing teeth pose a significantly more challenging task for automatic 3D landmark detection. This underscores the importance of evaluating landmark detection systems in scenarios that involve missing teeth. Finally, the CHaR-augmented version of PointMLP (CHaRNet) emerges as the best-performing method in all comparisons.
In terms of MEDE, CHaRNet achieves a remarkable 0.51 mm in typical dental models, 1.28 mm in all dentition models, and 3.94 mm when averaged across all dentition types. For MSR, it achieves an impressive 87.06% in typical dental models, 82.40% in all dentition models, and 68.95% when averaged across all dentition types.
@@ 5.2 Results
Table 5: Comparison of each neural network base model with its corresponding Conditioned Heatmap Regression (CHaR) version in terms of Mean Euclidean Distance Error (MEDE) expressed in millimeters (mm). The results are expressed across each dentition type and aggregated using the macro and micro average.
@@
00 0.82 0.86 1.26 0.74 0.77 0.59 0.81 0.70 0.65 0.51
01 3.37 2.20 3.66 2.66 2.66 1.86 2.64 2.11 2.29 1.74
02 4.52 3.25 5.38 3.33 5.54 2.04 3.74 3.04 4.32 3.09
03 7.83 5.68 7.81 4.47 4.57 6.90 3.24 2.98 6.01 6.73
04 4.26 1.93 9.32 5.15 6.87 2.79 4.25 5.71 4.39 3.76
10 1.09 1.04 1.10 0.87 0.99 0.82 0.89 0.84 0.88 0.62
11 4.43 4.37 5.00 3.34 3.29 2.49 4.24 3.04 2.59 1.75
12 6.21 4.56 7.96 5.95 6.45 6.83 6.00 4.69 4.95 4.89
13 7.65 7.60 8.85 8.97 10.55 2.66 2.92 4.46 1.51 4.83
14 17.13 21.54 17.64 16.39 8.57 12.44 15.53 14.47 15.37 11.54
macro-average 5.73 5.31 6.80 5.19 5.03 3.94 4.42 4.20 4.30 3.94
micro-average 1.99 1.71 2.46 1.67 1.90 1.40 1.72 1.46 1.57 1.28
Table 6: Comparison of each neural network base model with its corresponding Conditioned Heatmap Regression (CHaR) version in terms of Mean Success Rate (MSR) expresed in % of landmarks located withing a sphere of $1mm$ radius. The results are expressed across each dentition type and aggregated using the macro and micro average.
@@
00 68.58 68.16 75.32 77.46 77.63 83.38 70.12 73.70 78.35 87.06
01 52.71 54.90 62.92 66.04 70.83 76.67 60.52 63.44 70.00 79.48
02 54.33 57.08 59.00 66.00 60.50 76.42 59.00 65.75 63.33 78.42
03 33.75 36.88 43.75 52.50 64.38 65.63 54.38 55.63 56.88 53.75
04 37.50 48.75 27.50 36.25 36.25 60.00 41.25 31.25 47.50 48.75
10 60.00 60.83 72.03 72.20 69.76 76.42 62.84 66.72 70.57 80.47
11 39.79 44.38 50.63 56.04 58.33 66.46 45.63 53.33 61.88 73.33
12 46.75 50.88 52.38 60.38 56.25 64.88 47.25 56.38 64.13 73.25
13 35.00 38.75 37.50 50.00 20.00 65.00 50.00 48.75 63.75 66.25
14 18.75 13.75 17,50 26.25 50.00 40.00 16.25 23.75 41.25 48.75
macro-average 44.72 47.33 50.35 56.31 56.39 67.48 50.72 53.87 61.76 68.95
micro-average 61.46 62.03 69.31 72.12 71.52 78.71 64.29 68.48 72.97 82.40
Furthermore, qualitative results are shown in Figure 8 between the base and CHaR models, and the ground truth for different dental models. The red and green squares show wrong and correct landmark localization, respectively. This comparison clearly reveals superior performance from the CHaR models in complex dental models with missing teeth compared to their base model counterparts.
@@
Figure 8: Qualitative comparison between base models (top row), CHaR-augmented models (middle row) and ground truth (bottom row). Each column represents one of the point cloud learning models. The red and green squares represents wrong and correct landmark localizations, respectively.

### v5

@@ 5 Experiments
In this section, we show the effectiveness of CHaRM by coupling the CHaR module with several point cloud learning models spanning different families of algorithms and compare the best CHaR-based network (CHaRNet) with SOTA landmark detection models. The dataset splitting procedure is explained in Section 5.1. The experimental setup of each compared model is shown in 5.2. A description of the evaluation protocol used is introduced in Section 5.3. A comparison between baseline point cloud models and their corresponding CHaR variants is presented in Section 5.4. Finally, the comparison between CHaRNet and SOTA methods is introduced in Section 5.5.
5.1 Dataset splitting
We follow the typical three-way holdout setup for model comparison and performance estimation [62]. Specifically, the 1,214 dental models are divided into three subsets: training (70%), validation (15%), and testing (15%). To prevent patient-level correlation between subsets, we first group dental models by patient, ensuring that all models from a single patient belong exclusively to one of these subsets. Furthermore, to guaranty balanced representation of different anatomical complexities across these subsets, the splitting is conducted in a stratified manner according to the dental type taxonomy defined in Section 3.1. This stratification ensures that the distribution of dentition types remains consistent across training, validation, and testing subsets, enabling reliable evaluation of model performance on diverse dental anatomies.
5.2 Experimental setup
The experimental setup for each compared method in this work are as follows:
  • CHaR-based networks: The different CHaR-based versions were trained by minimizing the MSE loss (Formula 3) and the combined loss (Formula 5), respectively, using the Adam optimizer [63] for a total of 100 epochs and a batch size of 16. The initial learning rate is set to 0.005 and is progressively reduced using the cosine learning rate decay scheduler. To prevent overfitting, weight decay and dropout rate were set to 0.003 and 0.5, respectively, serving as regularization terms. With respect to hyper-parameters controlling the relative importance in the final loss value in the CHaR versions, $\lambda_{\text{reg}}$ and $\lambda_{\text{cls}}$ are set to 0.001 and 1 respectively, given the natural bigger quantity of the landmark localization loss.
  • TSMDL: Teeth segmentation network, iMeshSegNet, and landmark regression network, PointNet-Reg, are trained in IOSLandmarks-1k using default parameters.
  • ALIIOS: iMeshSegNet is used as teeth segmentation method since ALIIOS’ original teeth segmentation network, DentalModelSeg, is not available to be trained in IOSLandmarks-1k. We also change the default camera settings to accommodate the cameras’ positioning to our landmarks of interest described in Table 2.
All methods were trained in the same server with Intel Xeon E5-2698 as CPU and NVIDIA Tesla V100 32Gb as GPU.
5.3 Evaluation protocol
To rigorously evaluate landmark detection methods on dental models, especially when certain teeth (and thus their landmarks) might be absent, we evaluate the model in two clinically meaningful tasks: landmarks/tooth presence detection and landmark localization.
Landmarks/Tooth presence detection
We first evaluate the model’s ability to correctly determine the presence or absence of each tooth, which is equivalent to determining the presence or absence of each landmark since a set of five landmarks is associated with a tooth. Then, for every landmark prediction, four distinct outcomes are possible:
  • True Positive (TP): Landmark exists in the mesh and is correctly predicted in the mesh.
  • False Positive (FP): Landmark does not exist but is erroneously predicted in the mesh.
  • False Negative (FN): Landmark exists in the mesh but is erroneously predicted at the null point.
  • True Negative (TN): Landmark does not exist and is correctly predicted at the null point.
We compute the F1-score to evaluate landmarks/tooth classification performance:
$F1=\frac{2*TP}{2*TP+FP+FN}$ (7)
Landmark localization
For landmarks correctly identified as present (true positives), we measure how accurately the model predicts their 3D coordinates. Let $D_{i}=\|\hat{\mathbf{l}}_{i}-\mathbf{l}_{i}\|_{2}$ be the distance between the predicted and ground-truth landmark positions. To assess this accuracy, we employ two metrics:
  • Mean Euclidean Distance Error (MEDE): This metric calculates the average Euclidean distance between predicted and ground-truth landmark positions, computed exclusively over the true positives:
    $\text{MEDE}=\frac{1}{|TP|}\sum_{i\in TP}D_{i}$ (8)
  • Mean Success Ratio (MSR): To measure localization robustness, we compute the MSR at a clinically relevant threshold of 1 mm. MSR represents the percentage of true-positive landmark predictions falling within 1 mm of their ground-truth locations:
    $\text{MSR}=\frac{1}{|TP|}\sum_{i\in TP}\delta_{i}\times 100\%,\quad\delta_{i}=\begin{cases}1&\text{if }D_{i}\leq r,\\ 0&\text{if }D_{i}>r\end{cases}$ (9)
Motivation for the evaluation protocol
The motivation for this protocol arises from the clinical reality that landmark-presence detection and landmark localization are fundamentally different tasks. The presence detection reflects a classification problem (tooth existence), while localization assesses spatial accuracy once the existence is confirmed. Evaluating Euclidean distances involving the artificial "null point" (used to handle absent landmarks) would introduce arbitrary distance values with no clinical significance, thus introducing noise in the true localization performance evaluation. Our proposed protocol resolves this by evaluating localization accuracy strictly on landmarks correctly identified as present, ensuring both metrics remain clinically meaningful and transparent.
5.4 Base versus CHaR-based point cloud models
@@
Table 5: Comparison of each neural network base model with its corresponding Conditioned Heatmap Regression (CHaR) version in terms of F1-score for the task of teeth classification. The results are expressed across each dentition type and aggregated using the macro and micro averaged F1-score. Bold indicates the best F1-score per model.
@@
$0$ 0.95 0.87 0.87 0.99 0.93 0.99 0.93 0.99 0.93 0.99
01 0.84 0.98 0.84 0.97 0.85 0.92 0.85 0.86 0.91 0.98
02 0.93 0.96 0.79 0.95 0.80 0.97 0.87 0.95 0.81 0.95
03 0.89 0.93 0.87 0.93 0.91 0.89 0.91 0.91 0.89 0.85
04 0.87 0.93 0.68 0.87 0.87 0.93 0.81 0.81 0.81 0.93
10 0.99 0.99 0.99 0.99 0.97 0.99 0.97 0.99 0.99 0.99
11 0.94 0.95 0.93 0.95 0.97 0.97 0.94 0.96 0.95 0.98
12 0.91 0.94 0.91 0.91 0.90 0.93 0.91 0.93 0.93 0.92
13 0.81 0.87 0.75 0.75 0.75 0.93 0.81 0.87 0.87 0.90
14 0.56 0.68 0.56 0.62 0.62 0.81 0.62 0.68 0.62 0.68
Macro-avg 0.87 0.91 0.82 0.89 0.86 0.93 0.86 0.90 0.87 0.92
Micro-avg 0.97 0.98 0.96 0.98 0.95 0.98 0.98 0.98 0.98 0.99
We compare five base point cloud models (PointNet++, DGCNN, PointTransformer, PointNeXt, and PointMLP) against their CHaR-based versions along two complementary sources of error: (i) tooth/landmarks presence classification in Table 5, and (ii) landmark localization accuracy and robustness, only for the TP landmarks, in Table 6 for MEDE and Table 7 for MSR. Because only the CHaR-based versions have an explicit tooth presence output, for the base models we infer tooth presence heuristically: a tooth is considered present if all five of its associated landmarks are predicted inside the mesh; if any of them is sent to the null point, the tooth is treated as absent. This allows a fair, consistent definition of true positives for the localization metrics.
Table 5 shows that CHaR augmentation consistently improves the F1-score for tooth presence across architectures, with the largest gains in anatomically irregular cases (missing teeth / third molars). Better presence classification reduces the propagation of existence errors into localization, laying the foundation for more reliable downstream predictions.
When measuring localization performance, Tables 6 and 7 show that CHaR-based models generally improve spatial precision (lower MEDE) and increase the fraction of clinically acceptable predictions (higher MSR at 1 mm) across all point cloud models. The CHaR-based PointMLP model, which we call CHaRNet, exhibits the strongest localization profile: it achieves the lowest MEDE and highest MSR in the majority of dentition types, indicating both tighter errors and a larger fraction of landmarks falling within the clinically relevant 1 mm radius. For example, CHaRNet’s MSR improvement over its base counterpart is consistent across standard and challenging dentitions, reflecting robustness to anatomical variability. The reduction in MEDE demonstrates that, beyond simply detecting which landmarks to localize, CHaRNet refines their spatial estimates, producing more accurate coordinate predictions when the tooth is correctly recognized.
The improvements are especially meaningful in intermediate-to-hard cases where partial structure and missing teeth introduce uncertainty. In these settings, the conditioning provided by the CHaR module helps disambiguate the heatmap regression by weighting likelihoods according to tooth presence probabilities, effectively concentrating probability mass on the dentition. This leads to both fewer gross mislocalizations (reflected in MSR gains) and smaller average distance error (lower MEDE). That said, in extremely incomplete dentitions where the presence classification degrades, the conditioning can sometimes mislead localization, which is seen as occasional increases in MEDE for certain subclasses. This underscores that the efficacy of CHaRNet depends on reasonably clear teeth presence geometric signals.
Furthermore, qualitative results between the base network, CHaR-based network, and the ground truth for different point cloud methods and dental models are shown in Figure 8. The red and green squares show wrong and correct landmark localization, respectively. Base models struggle to detect/identify missing teeth on incomplete dental geometries by themselves. CHaR-based models tend to do a much better job across these complex anatomies.
In conclusion, CHaRM couples improved landmarks/tooth presence detection with more focused and reliable landmark regression. The addition of the CHaR module yields substantial gains in both precision and robustness, making it better suited for realistic clinical scenarios with variable and incomplete anatomy.
Table 6: Comparison of each neural network base model with its corresponding Conditioned Heatmap Regression (CHaR) version in terms of mean Euclidean distance error (MEDE) expressed in millimeters ($m$). The results are expressed across each dentition type and aggregated using the macro and micro averaged MEDE. Bold indicates the best MEDE per model.
@@
$0$ 0.90 0.90 0.83 0.82 0.75 0.65 0.85 0.80 0.73 0.56
01 1.47 1.40 1.34 1.22 0.88 0.87 1.10 1.15 1.10 0.76
02 1.33 1.44 1.51 1.51 1.29 0.99 1.30 1.25 1.18 0.99
03 2.68 1.91 2.04 2.20 0.89 1.15 1.30 1.38 1.89 1.66
04 3.33 3.26 3.54 3.48 4.77 2.71 2.68 2.76 3.85 2.14
10 0.98 0.97 0.82 0.85 0.83 0.73 0.91 0.86 0.82 0.64
11 2.04 1.84 1.83 1.83 1.12 1.12 1.76 1.53 0.97 0.98
12 1.47 1.52 1.67 1.41 0.97 0.84 1.28 1.14 0.83 0.67
13 2.27 1.31 1.20 0.95 3.31 0.92 1.27 1.18 0.98 0.85
14 4.54 3.60 5.21 4.61 1.42 1.92 4.58 3.56 2.75 1.91
Macro-avg 2.10 1.82 2.00 1.89 1.62 1.19 1.70 1.56 1.51 1.12
Micro-avg 1.15 1.08 1.02 1.01 0.88 0.76 0.99 0.95 0.85 0.70
Table 7: Comparison of each neural network base model with its corresponding Conditioned Heatmap Regression (CHaR) version in terms of mean success ratio (MSR) expresed in % of landmarks located withing a sphere of $1mm$ radius. The results are expressed across each dentition type and aggregated using the macro and micro averaged MSR. Bold indicates the best MSR per model.
Models
PointNet++ DGCNN PointTransformer PointNeXt PointMLP
Dent type Base CHaR Base CHaR Base CHaR Base CHaR Base CHaR
$0$ 63.9 64.4 74.0 74.2 74.8 81.1 66.1 73.7 75.2 85.2
01 47.3 47.1 62.9 62.9 70.2 73.8 57.0 63.4 67.1 77.6
02 50.0 52.3 61.3 62.8 63.7 73.4 55.0 65.7 64.1 79.5
03 25.7 36.2 45.0 40.7 69.4 69.6 45.0 55.6 43.0 49.1
04 6.6 11.1 11.1 24.4 9.3 40.0 15.5 31.2 20.0 29.5
10 58.6 59.4 72.0 71.4 69.2 75.8 61.9 66.7 69.9 79.7
11 41.7 40.0 52.9 54.5 59.2 64.5 46.0 53.3 62.2 70.9
12 51.7 49.7 60.9 65.7 63.8 72.9 50.5 56.3 69.5 80.1
13 21.0 38.3 46.5 60.0 30.1 61.6 39.6 48.7 55.0 69.0
14 20.0 31.1 33.3 35.5 61.2 66.6 18.0 23.7 48.8 63.8
Macro-avg 38.7 42.9 52.0 55.2 57.1 67.9 45.5 53.8 58.4 68.5
Micro-avg 55.5 57.8 66.7 68.9 69.2 75.8 60.6 68.4 70.1 79.8
@@
Figure 8: Qualitative comparison between base networks (top row), CHaR-based networks (middle row) and ground truth (bottom row). Each column represents one of the point cloud learning models. The red and green squares represents wrong and correct landmark localizations, respectively.
5.5 CHaRNet versus SOTA
We benchmark CHaRNet against the SOTA methods for IOS landmarks detection, TSMDL [8] and ALIIOS [9], using the evaluation method described in Section 5.3. For segmentation-based methods like TSMDL and ALIIOS, tooth presence classification is derived from the output of the teeth segmentation stage. Note that TSMDL and ALIIOS use the same teeth segmentation method, so results of tooth classification performance are the same. To ensure experimental fairness, all three methods were trained until convergence and evaluated using the same dataset split described in Section 5.1. The same hardware (CPU: Intel Core i7-13700H; GPU: NVIDIA RTX 4050) and the same batch size (1) were used to compute inference times in Table 10.
Table 8 shows the overall superior performance in the tooth classification task of CHaRNet versus TSMDL and ALIIOS. CHaRNet achieves the best F1-score across all dentition types, macro and micro averages, which further improves the qualitative results in the landmark localization task as shown in Figure 9.
Localization performance results across correctly classified teeth can be seen at Table 9. In terms of MEDE, CHaRNet beats TSMDL and ALIIOS across all dentition types, reducing the micro-averaged error to 0.80 mm. More importantly, CHaRNet achieves a macro-averaged error of 1.12 mm versus 1.90 mm and 2.17 mm from TSMDL and ALIIOS, respectively. When looking at MSR, we see that CHaRNet surpasses TSMDL and ALIIOS in all dentition types but in 00 and 10 subclasses, where TSMDL achieves slightly better results. These results show CHaRNet improves performance across complex dentitions with irregular geometries and missing teeth, which are understudied in previous works. This is exemplified in Figure 9: on standard dentitions (a) where the segmentation predictions are correct, ALIIOS and, especially, TSMDL deliver great results in landmark localization. However, in complex dentition cases (b), the segmentation usually fails and propagates this error to the individual tooth landmark detection models of ALIIOS and TSMDL. This failed segmentation does not allow a subsequent landmarks placement, since the individual tooth meshes are mostly incorrect. As can be seen, CHaRNet tends to behave better across standard and complex dentitions.
Table 10 compares the three methods in terms of time efficiency in CPU and GPU. CHaRNet is $4.5\times$ times faster in CPU and $14.8\times$ on GPU than TSMDL. ALIIOS is orders slower in both, especially in CPU. These big differences can be explained not only by the segmentation stage being time-consuming, but also by the necessity of TSMDL and ALIIOS to perform individual inferences for landmark localization for every tooth detected at the segmentation stage. ALIIOS is especially slower at individual tooth landmark localization given that five inferences per tooth must be performed.
In conclusion, CHaRNet achieves the best (i) tooth classification, (ii) landmark localization, and (iii) time efficiency at inference, allowing robust real-time landmark localization across any kind of dentition.
Table 8: Tooth/landmarks presence classification performance: F1-scores for ALIIOS, TSMDL, and CHaRNet. Results are shown for each dentition type, micro and macro-averaged. Bold indicates the best result for each dentition subclass.
Model $0$ 01 02 03 04 10 11 12 13 14 Macro-avg Micro-avg
ALIIOS 0.99 0.97 0.93 0.83 0.91 0.99 0.93 0.91 0.85 0.62 0.89 0.98
TSMDL 0.99 0.97 0.93 0.83 0.91 0.99 0.93 0.91 0.85 0.62 0.89 0.98
CHaRNet 0.99 0.98 0.95 0.85 0.93 0.99 0.98 0.92 0.90 0.68 0.92 0.99
Table 9: Landmarks localization performance: MEDE (↓) and MSR (↑) at 1 mm for ALIIOS, TSMDL, and CHaRNet. Results are shown for each dentition type, micro and macro-averaged. Bold indicates the best result for each dentition subclass.
Model $0$ 01 02 03 04 10 11 12 13 14 Macro-avg Micro-avg
MEDE (↓)
ALIIOS 0.93 1.49 1.64 3.15 4.39 0.94 1.76 1.56 2.69 3.10 2.17 1.12
TSMDL 0.59 1.26 1.30 2.75 2.85 0.64 1.48 1.51 2.31 3.27 1.80 0.83
CHaRNet 0.56 0.76 0.99 1.66 2.14 0.64 0.98 0.67 0.85 1.91 1.12 0.70
MSR (↑)
ALIIOS 70.2 55.9 55.0 22.7 12.5 70.5 53.9 58.2 41.1 40.0 48.0 65.7
TSMDL 87.6 71.3 70.3 36.4 15.5 85.1 67.3 65.9 52.7 37.7 59.0 79.4
CHaRNet 85.2 77.6 79.5 49.1 29.5 79.7 70.9 80.1 69.0 63.8 68.5 79.8
Table 10: Inference time performance: Mean execution time and standard deviation (in seconds) over the test set for ALIIOS, TSMDL, and CHaRNet. Bold denotes the best result for each device.
Model CPU (s) GPU (s)
ALIIOS 612.13 $\pm$ 24.27 24.70 $\pm$ 3.54
TSMDL 3.89 $\pm$ 0.85 3.55 $\pm$ 0.81
CHaRNet 0.86 $\pm$ 0.06 0.24 $\pm$ 0.03
(a) Standard dentition ($0$)
(b) Complex dentition (03)
Figure 9: Qualitative comparison between ALIIOS, TSMDL and CHaRNet on a standard dentition (a) and on a complex dentition (b). Predicition (top row) and ground-truth (bottom row) are shown for each model. Note that segmentation predictions and ground-truth are also provided for ALIIOS and TSMDL.